two equal charges of value q are placed at a distance of 2a and the third charge is at the midpoint  the potential energy of the system

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

Potential energy of the two ‘q’ charge system is\dfrac{k(q)(q)}{2a}
Potential energy of each of q and q’ charge system is\dfrac{k(q)(q')}{a}

Hence the total potential energy of the system is\dfrac{k(q)(q)}{2a}+2\dfrac{k(q)(q')}{a}